### v2

@@ 1 Introduction
The development of facial reaction generation has paralleled advances in deep learning. Early approaches using Convolutional Neural Networks (CNNs) established foundations in facial recognition and generation, such as Generative Adversarial Networks Goodfellow et al. (2014), Variational Autoencoders Kingma and Welling (2013), and flow-based models Dinh et al. (2016). The advent of Transformers Vaswani et al. (2017) enabled better processing of temporal sequences and multi-modal data, while diffusion models Song et al. (2020); Rombach et al. (2022) advanced the quality and variety of image generation.
@@
At the beginning, Luo et al. (2024); Song et al. (2019); Yoon et al. (2022) attempt to address these issues but typically focus on either uni-modal inputs or neglect the diversity of appropriate reactions. In applications, ignoring non-verbal inputs and repeating behaviors when interacting with users significantly impacts the realism of the agent Liang et al. (2023). This lack of diversity in facial reactions can result in a less engaging and natural interaction, further diminishing the agent’s perceived authenticity. Recent works have sought to improve the generation of multiple appropriate facial reactions by modeling a broader range of behaviors and contextual cues. For instance, BEAMER Hoque et al. (2023) introduces a Transformer-VAE architecture that leverages a behavioral encoder to process both the speaker’s behavior and texture. This encoder then maps this information into a latent space, which is used to generate a range of reactions from an appropriate listener. The framework emphasizes context-aware interaction, ensuring that the generated facial reactions are both contextually and emotionally aligned. Jun et al.  Yu et al. (2023) further enhance this by incorporating latent diffusion models. By modifying the traditional diffusion framework, they improve the model’s ability to capture and model the complex context of a conversation, enabling the generation of diverse and contextually relevant facial expressions. The inherent stochasticity of the diffusion process provides a natural mechanism for the generation of multiple plausible reactions, which enhances the realism and flexibility of the agent’s responses. In a similar vein, Nguyen et al.  Nguyen et al. (2024) introduce a novel framework leveraging vector quantized diffusion models for generating multiple appropriate facial reactions in dyadic interactions. Liu et al.  Liu et al. (2024) tackle the challenge of generating a range of appropriate reactions using discrete latent variables. Their method explicitly models the one-to-many mapping problem, ensuring that multiple, contextually appropriate facial reactions can be generated from the same initial input, thereby addressing the issue of overfitting to a single reaction type. Finally, a recent approach by Zhu et al.  Zhu et al. (2024) introduces a personalized approach to this problem. By incorporating weight-editing mechanisms into the generation process, their model allows for fine-tuning of facial reactions based on personalized data, such as user-specific preferences or emotional cues. This approach offers a more tailored and individualized set of reactions, further improving the agent’s capacity to engage with users in a more realistic and empathetic manner.
@@
The REACT2023 Song et al. (2023) and REACT2024 Song et al. (2024) challenges formalize these requirements by introducing the Multiple Appropriate Reaction Generation problem, which emphasizes the need to generate various appropriate reactions to the same context while maintaining realism. Specifically, given a speaker’s dyadic interaction video and audio, the model generates multiple sequences of the listener’s facial reactions, comprising 3D Morphable Face Model coefficients, Facial Action Units, Valence and Arousal, and Facial Expression.
@@
We evaluate ReactDiff on the official dataset of the REACT2024 challenge Song et al. (2024), which is constructed from the existing RECOLA Ringeval et al. (2013) and NoXI Cafaro et al. (2017) datasets. Experiments demonstrate that ReactDiff can generate more realistic, appropriate, and diverse facial reactions. The main contributions of this work are as follows:
@@ 2.1 Facial Reaction Generation
The study of facial reaction generation emerged with the development of Human-Computer Interaction in the early 1990s. Takeuchi et al. introduced facial displays as a new modality in computer-human interaction Takeuchi and Nagao (1993), while Poggi et al. established fundamental theories of non-verbal communication systems, focusing on gaze and facial expressions Poggi and Pelachaud (2000). Early computational approaches were developed by Ochs et al., who computed facial expressions using a Facial Expression of Emotions Blending Module Ochs et al. (2005), and Zeng et al., who examined methods for understanding human affective behaviour in real-world settings Zeng et al. (2007).
With the advent of deep learning, more sophisticated approaches for facial reaction generation have been proposed. DyadGan Huang and Khan (2017) and StlyeFaceV Qiu et al. (2022) pioneered the use of generative models for facial expressions in dyadic interactions. More recent frameworks like ReactFace Luo et al. (2024) and UniFaRN Liang et al. (2023) have further advanced the field by addressing the challenges of balancing appropriateness, realism, and diversity. Chen et al. propose a diverse graph transformers model that incorporates both spatial dual-graphs and temporal hyperbolic-graph to capture expression-related structures from facial videos Chen et al. (2024).
A significant breakthrough in facial reaction generation came with the development of 3D Morphable Face Models (3DMM) Blanz and Vetter (1999). These statistical models have become powerful tools for modeling facial geometry and appearance by capturing facial variability across populations. Recent studies have bridged the gap between 2D and 3D representations by proposing methods to generate 2D facial image sequences from 3D facial behavioursWang et al. (2022); Ren et al. (2021), enabling high-quality facial reaction generation through the integration of 2D facial landmarks and 3D facial geometry.
@@ 2.2 Multi-Modal Fusion
Transformer, proposed by Vaswani et al. in 2017 Vaswani et al. (2017), has revolutionized deep learning. Initially designed for natural language processing, its self-attention mechanism allows for efficient modeling of long-range dependencies. The architecture’s success quickly spreads to other domains: Vision Transformer (ViT)Dosovitskiy (2020) adapts for image processing, while models like CLIPRadford et al. (2021) and GPTRadford (2018) demonstrate effectiveness in multi-modal tasks and large-scale language modeling. More recently, Diffusion Transformers (DiT) Peebles and Xie (2023) have shown promise in generation tasks, further expanding the architecture’s versatility across various AI applications. Building upon the success of Transformer architecture, numerous approaches have been developed to handle multi-modal fusion tasks. The BERT Devlin et al. (2018) architecture is extended to a multi-modal two-stream model, processing visual and textual inputs in separate interacting streams Lu et al. (2019). Significant progress has also been made in learning transferable representations across modalities. Radford et al.Radford et al. (2021) introduced CLIP, learning transferable visual representations from natural language supervision. BLIPLi et al. (2022) further advances this field by employing a bootstrapping approach for vision-language understanding and generation.
Recent research has explored emotion understanding in multi-modal contexts. Kumar et.al Kumar and Raman (2022) proposed a BERT-based dual-channel system for text emotion recognition that demonstrates the effectiveness of transformer-based architectures in emotional understanding tasks. To address the challenge of incomplete data, Cheng et al. propose a transformer autoencoder that leverages both intra-modality and inter-modality transformers Cheng et al. (2024). In a similar vein, Sadok et al. proposed a multimodal dynamical variational autoencoder that effectively fuses audiovisual features through disentangled latent representations Sadok et al. (2024).
@@ 2.3 Diffusion Models
Diffusion Models (DMs) represent a significant advancement in image generation. Diffusion Probabilistic Models (DPMs) Ho et al. (2020) consist of a forward process (mapping signal to noise) and a reverse process (mapping noise to signal). They can denoise an arbitrary Gaussian noise map to a clean image sample after successive denoising passes proposed learning a function to predict noise using a UNet. To make DPMs more practical, Denoising Diffusion Implicit Models (DDIMs) Song et al. (2020) first propose a method of implicit sampling through DPMs and accelerated the sampling speed. To address computational challenges, Rombach et al. applied DMs in the latent space of pre-trained autoencoders Rombach et al. (2022). For face image processing, Diffusion Autoencoders (DiffAE) Preechakul et al. (2022) encode facial images into two-part latent codes, using a DPM as the decoder for stochastic variations. Building on this, Diffusion Video Autoencoder Kim et al. (2023) is proposed for human face video editing, offering superior reconstruction performance and disentanglement of identity features. Latent Diffusion Models (LDMs) have evolved significantly since their inception. Early LDMs, such as Stable Diffusion 1.5 XL, primarily rely on CNN-based architectures Rombach et al. (2022). However, recent advancements have seen a shift towards pure Transformer structures, as exemplified by the Diffusion Transformers (DiT) model Peebles and Xie (2023). Furthermore, LDMs have demonstrated the ability to be controlled by multiple modal conditions, allowing for more diverse and specific generation outcomes. For instance, techniques like ControlNet Zhang et al. (2023) have enabled fine-grained control over the generation process using various types of conditioning information.
@@ 3.1 Technical Overview and Comparison with Existing Methods
Traditional approaches to facial reaction generation have followed two main paradigms: transformer-based methods that excel at temporal modeling but struggle with diversity, and diffusion-based methods that enable diverse generation but lack fine-grained control over multi-modal interactions. For example, BEAMER Hoque et al. (2023) uses a pure transformer architecture that achieves good temporal consistency but produces limited variation in reactions. Conversely, PerFRDiff Zhu et al. (2024) employs diffusion models that generate diverse outputs but may not maintain consistent temporal relationships between audio and visual features. ReactDiff advances beyond these approaches through three key innovations. First, while existing methods typically process audio and visual features independently before late fusion, our Multi-Modality Transformer (MMT) employs a hierarchical attention structure with both intra- and inter-cross attention, enabling more nuanced multi-modal interactions. Second, unlike previous work that applies diffusion models directly to the output space, we incorporate the diffusion process in the latent space between MMT’s encoder and decoder, allowing for better balance between diversity and temporal consistency. Third, our conditional DDIM uses the speaker’s behavior constraint vector to guide the generation process, ensuring that diverse reactions remain contextually appropriate. This unique architecture allows ReactDiff to maintain the temporal modeling capabilities of transformers while leveraging the generative power of diffusion models, addressing the limitations of previous approaches.
The challenge of generating diverse and realistic facial reactions necessitates advanced modeling techniques. While Variational Autoencoders (VAEs) and Generative Adversarial Networks (GANs) are initially considered, they present limitations in balancing information retention, sampling quality, and training stability. Diffusion Probabilistic Models (DPMs) excel in stability but struggle with multi-modal and temporal information processing Preechakul et al. (2022). To address these limitations, we propose the Facial Reaction Generation Diffusion Model (ReactDiff), integrating cross-attention mechanisms and Transformer layers. ReactDiff comprises three primary modules: feature extraction, Multi-Modality Transformer (MMT), and latent diffusion model. As illustrated in Fig 2(a), the framework first extracts facial and acoustic features from raw input videos. The MMT module then reconstructs the listener’s reaction, including emotional states and 3D Morphable Model (3DMM) coefficients. Finally, in the latent reaction space, we process the MMT-produced latent vectors to enhance diversity and realism of generated reactions, constrained by corresponding speaker information. In summary, our design enables complex modeling of relationships within and across modalities.
@@ 3.2.1 Facial features
For facial reaction generation, raw videos often contain task-irrelevant pixels like background and clothing. When trained with limited samples, the model has the risk of being over-fitted (commonly called shortcut learning Geirhos et al. (2020)) to these irrelevant pixels, ignoring reaction-related features such as facial movements.
@@
  • The AUs based on the Facial Action Coding System (FACS) Ekman and Friesen (1978) describe the fundamental actions of individuals or groups of muscles on the face. Each AU corresponds to a specific muscular activity that produces a distinct facial movement. For example, AU12 represents the contraction of the zygomatic major muscle, which produces a smile. By controlling the intensity and combination of different AUs, a wide range of facial expressions can be generated.
@@
  • ELIM Kim and Song (2022) is an identity matching method using the Sinkhorn-Knopp algorithm, where reference and target IDs’ feature sets are encoded, and optimal matching flows are generated to produce ID-invariant emotional representations. We extract VA features $f^{VA}$ by ELIM encoder.
  • ME-GraphAU Luo et al. (2022) is a method for modeling Action Unit relationships using a combination of node feature learning and multi-dimensional edge feature learning to create an AU relation graph. The AU features $f^{AU}$ are extracted from ME-GraphAU without a classifier.
  • ResMaskNet Pham et al. (2021) uses several Masking Blocks, which are applied across Residual Layers to improve the network’s attention ability on important information. The FE features $f^{FE}$ are extracted from ResMaskNet without a classifier.
@@ 3.2.2 Acoustic features
We also preprocess the audio for the network. Wav2vec 2.0 Baevski et al. (2020) is a self-supervised framework for learning speech representations directly from raw audio data, achieving high performance on speech recognition tasks with minimal labeled data. Mel Frequency Cepstral Coefficients (MFCC) features are derived from the Fourier transform of a signal, representing the short-term power spectrum of sound on a mel scale, commonly used in speech and audio processing. The audio is extracted as MFCC features and Audio Feature Representations (AFR). The MFCC and AFR are aligned with video frames, as in Equation 2.

### v3

@@ 1 Introduction
The development of facial reaction generation has paralleled advances in deep learning. Early approaches using Convolutional Neural Networks (CNNs) established foundations in facial recognition and generation, such as Generative Adversarial Networks (Goodfellow et al., 2014), Variational Autoencoders (Kingma and Welling, 2013), and flow-based models (Dinh et al., 2016). The advent of Transformers (Vaswani et al., 2017) enabled better processing of temporal sequences and multi-modal data, while diffusion models (Song et al., 2020; Rombach et al., 2022) advanced the quality and variety of image generation.
@@
At the beginning, (Luo et al., 2024; Song et al., 2019; Yoon et al., 2022) attempt to address these issues but typically focus on either uni-modal inputs or neglect the diversity of appropriate reactions. In applications, ignoring non-verbal inputs and repeating behaviors when interacting with users significantly impacts the realism of the agent (Liang et al., 2023). This lack of diversity in facial reactions can result in a less engaging and natural interaction, further diminishing the agent’s perceived authenticity. Recent works have sought to improve the generation of multiple appropriate facial reactions by modeling a broader range of behaviors and contextual cues. For instance, BEAMER (Hoque et al., 2023) introduces a Transformer-VAE architecture that leverages a behavioral encoder to process both the speaker’s behavior and texture. This encoder then maps this information into a latent space, which is used to generate a range of reactions from an appropriate listener. The framework emphasizes context-aware interaction, ensuring that the generated facial reactions are both contextually and emotionally aligned. Jun et al.  (Yu et al., 2023) further enhance this by incorporating latent diffusion models. By modifying the traditional diffusion framework, they improve the model’s ability to capture and model the complex context of a conversation, enabling the generation of diverse and contextually relevant facial expressions. The inherent stochasticity of the diffusion process provides a natural mechanism for the generation of multiple plausible reactions, which enhances the realism and flexibility of the agent’s responses. In a similar vein, Nguyen et al.  (Nguyen et al., 2024) introduce a novel framework leveraging vector quantized diffusion models for generating multiple appropriate facial reactions in dyadic interactions. Liu et al.  (Liu et al., 2024) tackle the challenge of generating a range of appropriate reactions using discrete latent variables. Their method explicitly models the one-to-many mapping problem, ensuring that multiple, contextually appropriate facial reactions can be generated from the same initial input, thereby addressing the issue of overfitting to a single reaction type. Finally, a recent approach by Zhu et al.  (Zhu et al., 2024) introduces a personalized approach to this problem. By incorporating weight-editing mechanisms into the generation process, their model allows for fine-tuning of facial reactions based on personalized data, such as user-specific preferences or emotional cues. This approach offers a more tailored and individualized set of reactions, further improving the agent’s capacity to engage with users in a more realistic and empathetic manner.
@@
The REACT2023 (Song et al., 2023) and REACT2024 (Song et al., 2024) challenges formalize these requirements by introducing the Multiple Appropriate Reaction Generation problem, which emphasizes the need to generate various appropriate reactions to the same context while maintaining realism. Specifically, given a speaker’s dyadic interaction video and audio, the model generates multiple sequences of the listener’s facial reactions, comprising 3D Morphable Face Model coefficients, Facial Action Units, Valence and Arousal, and Facial Expression.
@@
We evaluate ReactDiff on the official dataset of the REACT2024 challenge (Song et al., 2024), which is constructed from the existing RECOLA (Ringeval et al., 2013) and NoXI (Cafaro et al., 2017) datasets. Experiments demonstrate that ReactDiff can generate more realistic, appropriate, and diverse facial reactions. The main contributions of this work are as follows:
@@ 2.1 Facial Reaction Generation
The study of facial reaction generation emerged with the development of Human-Computer Interaction in the early 1990s. Takeuchi et al. introduced facial displays as a new modality in computer-human interaction (Takeuchi and Nagao, 1993), while Poggi et al. established fundamental theories of non-verbal communication systems, focusing on gaze and facial expressions (Poggi and Pelachaud, 2000). Early computational approaches were developed by Ochs et al., who computed facial expressions using a Facial Expression of Emotions Blending Module (Ochs et al., 2005), and Zeng et al., who examined methods for understanding human affective behaviour in real-world settings (Zeng et al., 2007).
With the advent of deep learning, more sophisticated approaches for facial reaction generation have been proposed. DyadGan (Huang and Khan, 2017) and StlyeFaceV (Qiu et al., 2022) pioneered the use of generative models for facial expressions in dyadic interactions. More recent frameworks like ReactFace (Luo et al., 2024) and UniFaRN (Liang et al., 2023) have further advanced the field by addressing the challenges of balancing appropriateness, realism, and diversity. Chen et al. propose a diverse graph transformers model that incorporates both spatial dual-graphs and temporal hyperbolic-graph to capture expression-related structures from facial videos (Chen et al., 2024).
A significant breakthrough in facial reaction generation came with the development of 3D Morphable Face Models (3DMM) (Blanz and Vetter, 1999). These statistical models have become powerful tools for modeling facial geometry and appearance by capturing facial variability across populations. Recent studies have bridged the gap between 2D and 3D representations by proposing methods to generate 2D facial image sequences from 3D facial behaviours(Wang et al., 2022; Ren et al., 2021), enabling high-quality facial reaction generation through the integration of 2D facial landmarks and 3D facial geometry.
@@ 2.2 Multi-Modal Fusion
Transformer, proposed by Vaswani et al. in 2017 (Vaswani et al., 2017), has revolutionized deep learning. Initially designed for natural language processing, its self-attention mechanism allows for efficient modeling of long-range dependencies. The architecture’s success quickly spreads to other domains: Vision Transformer (ViT)(Dosovitskiy, 2020) adapts for image processing, while models like CLIP(Radford et al., 2021) and GPT(Radford, 2018) demonstrate effectiveness in multi-modal tasks and large-scale language modeling. More recently, Diffusion Transformers (DiT) (Peebles and Xie, 2023) have shown promise in generation tasks, further expanding the architecture’s versatility across various AI applications. Building upon the success of Transformer architecture, numerous approaches have been developed to handle multi-modal fusion tasks. The BERT (Devlin et al., 2018) architecture is extended to a multi-modal two-stream model, processing visual and textual inputs in separate interacting streams (Lu et al., 2019). Significant progress has also been made in learning transferable representations across modalities. Radford et al.(Radford et al., 2021) introduced CLIP, learning transferable visual representations from natural language supervision. BLIP(Li et al., 2022) further advances this field by employing a bootstrapping approach for vision-language understanding and generation.
Recent research has explored emotion understanding in multi-modal contexts. Kumar et.al (Kumar and Raman, 2022) proposed a BERT-based dual-channel system for text emotion recognition that demonstrates the effectiveness of transformer-based architectures in emotional understanding tasks. To address the challenge of incomplete data, Cheng et al. propose a transformer autoencoder that leverages both intra-modality and inter-modality transformers (Cheng et al., 2024). In a similar vein, Sadok et al. proposed a multimodal dynamical variational autoencoder that effectively fuses audiovisual features through disentangled latent representations (Sadok et al., 2024).
@@ 2.3 Diffusion Models
Diffusion Models (DMs) represent a significant advancement in image generation. Diffusion Probabilistic Models (DPMs) (Ho et al., 2020) consist of a forward process (mapping signal to noise) and a reverse process (mapping noise to signal). They can denoise an arbitrary Gaussian noise map to a clean image sample after successive denoising passes proposed learning a function to predict noise using a UNet. To make DPMs more practical, Denoising Diffusion Implicit Models (DDIMs) (Song et al., 2020) first propose a method of implicit sampling through DPMs and accelerated the sampling speed. To address computational challenges, Rombach et al. applied DMs in the latent space of pre-trained autoencoders (Rombach et al., 2022). For face image processing, Diffusion Autoencoders (DiffAE) (Preechakul et al., 2022) encode facial images into two-part latent codes, using a DPM as the decoder for stochastic variations. Building on this, Diffusion Video Autoencoder (Kim et al., 2023) is proposed for human face video editing, offering superior reconstruction performance and disentanglement of identity features. Latent Diffusion Models (LDMs) have evolved significantly since their inception. Early LDMs, such as Stable Diffusion 1.5 XL, primarily rely on CNN-based architectures (Rombach et al., 2022). However, recent advancements have seen a shift towards pure Transformer structures, as exemplified by the Diffusion Transformers (DiT) model (Peebles and Xie, 2023). Furthermore, LDMs have demonstrated the ability to be controlled by multiple modal conditions, allowing for more diverse and specific generation outcomes. For instance, techniques like ControlNet (Zhang et al., 2023) have enabled fine-grained control over the generation process using various types of conditioning information.
@@ 3.1 Technical Overview and Comparison with Existing Methods
Traditional approaches to facial reaction generation have followed two main paradigms: transformer-based methods that excel at temporal modeling but struggle with diversity, and diffusion-based methods that enable diverse generation but lack fine-grained control over multi-modal interactions. For example, BEAMER (Hoque et al., 2023) uses a pure transformer architecture that achieves good temporal consistency but produces limited variation in reactions. Conversely, PerFRDiff (Zhu et al., 2024) employs diffusion models that generate diverse outputs but may not maintain consistent temporal relationships between audio and visual features. ReactDiff advances beyond these approaches through three key innovations. First, while existing methods typically process audio and visual features independently before late fusion, our Multi-Modality Transformer (MMT) employs a hierarchical attention structure with both intra- and inter-cross attention, enabling more nuanced multi-modal interactions. Second, unlike previous work that applies diffusion models directly to the output space, we incorporate the diffusion process in the latent space between MMT’s encoder and decoder, allowing for better balance between diversity and temporal consistency. Third, our conditional DDIM uses the speaker’s behavior constraint vector to guide the generation process, ensuring that diverse reactions remain contextually appropriate. This unique architecture allows ReactDiff to maintain the temporal modeling capabilities of transformers while leveraging the generative power of diffusion models, addressing the limitations of previous approaches.
The challenge of generating diverse and realistic facial reactions necessitates advanced modeling techniques. While Variational Autoencoders (VAEs) and Generative Adversarial Networks (GANs) are initially considered, they present limitations in balancing information retention, sampling quality, and training stability. Diffusion Probabilistic Models (DPMs) excel in stability but struggle with multi-modal and temporal information processing (Preechakul et al., 2022). To address these limitations, we propose the Facial Reaction Generation Diffusion Model (ReactDiff), integrating cross-attention mechanisms and Transformer layers. ReactDiff comprises three primary modules: feature extraction, Multi-Modality Transformer (MMT), and latent diffusion model. As illustrated in Fig 2(a), the framework first extracts facial and acoustic features from raw input videos. The MMT module then reconstructs the listener’s reaction, including emotional states and 3D Morphable Model (3DMM) coefficients. Finally, in the latent reaction space, we process the MMT-produced latent vectors to enhance diversity and realism of generated reactions, constrained by corresponding speaker information. In summary, our design enables complex modeling of relationships within and across modalities.
@@ 3.2.1 Facial features
For facial reaction generation, raw videos often contain task-irrelevant pixels like background and clothing. When trained with limited samples, the model has the risk of being over-fitted (commonly called shortcut learning (Geirhos et al., 2020)) to these irrelevant pixels, ignoring reaction-related features such as facial movements.
@@
  • The AUs based on the Facial Action Coding System (FACS) (Ekman and Friesen, 1978) describe the fundamental actions of individuals or groups of muscles on the face. Each AU corresponds to a specific muscular activity that produces a distinct facial movement. For example, AU12 represents the contraction of the zygomatic major muscle, which produces a smile. By controlling the intensity and combination of different AUs, a wide range of facial expressions can be generated.
@@
  • ELIM (Kim and Song, 2022) is an identity matching method using the Sinkhorn-Knopp algorithm, where reference and target IDs’ feature sets are encoded, and optimal matching flows are generated to produce ID-invariant emotional representations. We extract VA features $f^{VA}$ by ELIM encoder.
  • ME-GraphAU (Luo et al., 2022) is a method for modeling Action Unit relationships using a combination of node feature learning and multi-dimensional edge feature learning to create an AU relation graph. The AU features $f^{AU}$ are extracted from ME-GraphAU without a classifier.
  • ResMaskNet (Pham et al., 2021) uses several Masking Blocks, which are applied across Residual Layers to improve the network’s attention ability on important information. The FE features $f^{FE}$ are extracted from ResMaskNet without a classifier.
@@ 3.2.2 Acoustic features
We also preprocess the audio for the network. Wav2vec 2.0 (Baevski et al., 2020) is a self-supervised framework for learning speech representations directly from raw audio data, achieving high performance on speech recognition tasks with minimal labeled data. Mel Frequency Cepstral Coefficients (MFCC) features are derived from the Fourier transform of a signal, representing the short-term power spectrum of sound on a mel scale, commonly used in speech and audio processing. The audio is extracted as MFCC features and Audio Feature Representations (AFR). The MFCC and AFR are aligned with video frames, as in Equation 2.